### v1

@@ Abstract
Cardiovascular disease has become increasingly prevalent in modern society and has a significant effect on global health and well-being. Heart-related conditions are intricate, multifaceted disorders, which may be influenced by a combination of genetic predispositions, lifestyle choices, and various socioeconomic and clinical factors. Information regarding these potentially complex interrelationships is dispersed among diverse types of textual data, which include patient narratives, medical records, and scientific literature, among others. Natural language processing (NLP) techniques have increasingly been adopted as a powerful means to analyse and make sense of this vast amount of unstructured data. This, in turn, can allow healthcare professionals to gain deeper insights into the cardiology field, which has the potential to revolutionize current approaches to the diagnosis, treatment, and prevention of cardiac problems. This review provides a detailed overview of NLP research in cardiology between 2014 and 2025. We queried six literature databases to find articles describing the application of NLP techniques in the context of a range of different cardiovascular diseases. Following a rigorous screening process, we identified a total of 265 relevant articles. We analysed each article from multiple dimensions, i.e., NLP paradigm types, cardiology-related task types, cardiovascular disease types, and data source types. Our analysis reveals considerable diversity within each of these dimensions, thus demonstrating the considerable breadth of NLP research within the field. We also perform a temporal analysis, which illustrates the evolution and changing trends in NLP methods employed over the last decade that we cover. To our knowledge, the review constitutes the most comprehensive overview of NLP research in cardiology to date.
@@ 1 Introduction
Cardiovascular diseases (CVDs) comprise a group of disorders that affect the heart and blood vessels, posing a major global health challenge. The primary types include coronary artery disease, heart failure, arrhythmias, valvular heart disease, and hypertension. These conditions can significantly diminish an individual’s quality of life and overall well-being, with their prevalence continuing to rise due to aging populations and lifestyle changes [1]. According to the World Health Organization, CVDs are the leading cause of death worldwide, accounting for an estimated 17.9 million deaths each year [2]. Consequently, early detection and effective management of CVDs are essential to improving patient outcomes and mitigating their burden on healthcare systems.
@@ 2.1 Search strategy
Table 1: Keywords for Literature Search
@@ 3 Findings
  • The type of NLP method employed;
  • The cardiology-related task/application addressed by the research;
  • The type of cardiovascular disease that forms the focus of the research;
  • The type of dataset to which the reported NLP method was applied.
@@ 3.1 Data Sources
Figure 3: Hierarchical Dataset Classification Schema.
@@ 3.1.1 Top-level Category
Figure 4: Treemap of datasets.
@@
As shown in figure 4, a total of 9 different types of textual data have been considered in the reviewed articles. The vast majority of studies make use of Electronic Health Records (EHRs), which account for 82.6% of the data sources used in the studies included in the review. Published literature and medical databases constitute the second largest category at 4.7%, followed closely by synthetic or generated data at 4.3%. Clinical Trial Data represents 2.7% of the sources, while telehealth data accounts for 1.9%. Patient-reported outcomes and surveys make up 1.6%, and home healthcare data represents 1.2%. Administrative data (0.8%) and social media data (0.4%) represent smaller portions of the data sources used.
@@
Synthetic or generated data.
@@
Administrative data
Administrative data encompasses structured information generated during healthcare business operations, including insurance claims, billing records, and operational metrics, which is typically maintained in systems separate from but complementary to Electronic Health Records (EHRs, [27]). While primarily collected for billing, reimbursement, and administrative purposes, source allocation, and system-level performance indicators across different healthcare settings, administrative data has become a valuable resource for healthcare research due to its comprehensiveness and longitudinal nature. In the reviewed papers, Sharma [28] utilised administrative health data to predict the risk of 30-day readmissions in patients with heart failure.
@@
Clinical trial data represents highly structured and carefully collected information from controlled research studies investigating medical interventions, treatments, or diagnostic approaches [29]. This data type is characterized by its rigorous collection protocols, standardized formats, and detailed documentation of patient outcomes and adverse events. Clinical trial data is considered the gold standard for evidence-based medicine, though its generalizability may be limited by strict inclusion/exclusion criteria. For example, Cunningham et al. [30] utilized the INVESTED trial data [31] to validate and fine-tune a Natural Language Processing (NLP) model for automatically identifying heart failure hospitalizations.
Home Healthcare Data
Home healthcare data encompasses clinical information collected during healthcare services provided in patients’ homes, including vital signs, symptoms, medication adherence, and functional assessments [32]. This data type offers unique insights into patients’ daily living conditions and health management behaviours outside traditional clinical settings. Home healthcare data is particularly valuable for understanding chronic disease management and patient self-care capabilities. For example, Chae [33] used 2.3 million home healthcare clinical notes to identify patients in home healthcare with heart failure symptoms and poor self-management (SM).
@@
Patient-reported Outcomes and Surveys
@@ 3.1.3 Clinical Documentation and Diagnostic Information
Figure 6: Bar Chart of sub-subcategories of Clinical Documentation and Diagnostic Information in EHRs.
@@ 3.1.4 Imaging and Radiology Reports
Figure 7: Bar Chart of different cardiac imaging and radiology reports
@@
Figure 8: Co-occurrence of different cardiac imaging and radiology reports
The co-occurrence map, as shown in figure 8, reveals several significant patterns in how these different report types are used together in research. Echocardiogram reports frequently co-occur with Chest X-ray reports (n=4) and ECG reports (n=4), suggesting common research approaches combining anatomical imaging with functional assessment. Cardiac MRI and CT reports show co-occurrence with each other (n=1), indicating complementary use of these advanced imaging techniques. Coronary Angiography reports demonstrate integration with multiple modalities, co-occurring with Echocardiogram reports (n=1), ECG reports (n=1), and Cardiac Catheterization reports (n=1), reflecting the comprehensive approach often needed in cardiovascular research.
@@ 3.3 Target Cardiovascular Diseases
In Coronary Artery Disease research, NLP has been particularly valuable in analyzing cardiac catheterization reports [94, 95, 96] and identifying acute coronary events from clinical documentation [4, 97, 98]. Studies in this category have also focused on developing predictive models for disease onset and outcome assessment [99].
@@ 3.4.1 Rule-Based Methods
Rule-based methods use predefined patterns, dictionaries, and expert-crafted rules to extract information from text. These methods have proven particularly effective for extracting standardized measurements and well-defined clinical entities. For example, Wagholikar [111] extracted ejection fraction values from echocardiogram reports, which achieved high precision due to the consistent formatting of these measurements. Ambrosy [21] utilized rule-based approaches to leverage standardized clinical criteria for identifying worsening heart failure events. Olivella et al. [112] used rule-based NLP approaches to filter cases of heart failure patients. Mefford et al. [113] used rule-based NLP methods to examine variation in worsening heart failure hospitalizations, considering multiple factors. There are also other works using rule-based statistics (e.g., Bag-of-Words, Skip-gram) as features for machine learning algorithms, successfully applied in scenarios such as random forest [114, 115] and support vector machine [116, 84, 117].
@@ 3.4.3 Deep Learning-based Methods
Since the release of ChatGPT in 2022, Large Language Models (LLMs) have marked a major paradigm shift in NLP. LLMs are huge Transformer-based models (usually billions of parameters) pre-trained on massive scale of web data (usually trillions of tokens) to obtain outstanding general problem-solving capabilities and strong in-context learning capabilities, which enable LLMs to adapt to new tasks without modifying the parameters. LLMs have become an emerging dominant paradigm in cardiology, significantly outperforming previous methods in risk prediction for cardiovascular diseases [180, 81, 149], heart failure [182], and automated echocardiography report analysis [181]. Many works also explored utilizing LLMs for assisting diagnosis of related diseases, including as knowledge sources [199], agent for symptom description and management [200], and interpreters of existing reports [201]. LLMs also show generalized ability to handle signals from multiple sources, including ECG signals [70, 202, 203], text-based cardiology disease reports [204, 205, 206], and educational materials [74, 86, 207, 208]. Due to their auto-regressive nature, LLMs are also widely used in text generation tasks, successfully applied in question answering for cardiology-related topics [24, 62, 183, 184], AI assistants for clinical advice on heart attack [185, 186], atrial fibrillation [77, 187], cardiovascular rehabilitation [188], and cardiac electrophysiology procedures [209, 7], and writing summaries of cardiology reports [65], cardiovascular magnetic resonance reports [189, 210], and education materials [190], etc. LLMs are also explored as research and decision-support tools for domain experts [191, 211, 212]. In practical scenarios, the high-quality LLM-generated content can significantly lower the burden of domain experts and the cost of patients for professional advice. Another application focuses on extracting information from texts via prompt-based instructions. For example, Xu et al. [71] leveraged LLMs to extract structured knowledge from heart failure-related texts and built the first knowledge graph in this domain. Shang et al. [192] evaluated the factuality of extracted Cardiovascular-Kidney-Metabolic syndrome knowledge within LLMs to ensure their reliability. Chi et al. [193] utilized LLMs to extract echocardiogram entities in a flexible manner. Tully et al. [194] compared LLMs to rank the priorities for surgical patients needing preoperative cardiac evaluations. Vasisht Shankar et al. [213] converted text-based cardiovascular diagnostic reports into structured datasets with LLMs. Finally, due to LLMs’ high hallucination rate in knowledge-intensive scenarios, the retrieval-augmented generation (RAG) technique has been explored in cardiology, which aims to enhance LLM-generated content with retrieved knowledge or texts. For example, Adejumo et al. [80] incorporated structured information from inpatient heart failure notes to combine into the context of LLMs. Hayama et al. [195] comprehensively evaluated RAG-based LLMs and general LLMs in Cardiology specialist examinations in Japan, proving external knowledge useful for QA tasks. Lakhdhar et al. [196] used RAG in developing AI assistants for personalized cardiovascular risk assessment, and Parameswaran et al. [197] further proved its effectiveness by using RAG to enhance LLM performance in providing nutrition information for cardiovascular diseases.
@@ 4.1 Main Challenges
Data Privacy and Compliance with Regulation
@@ 4.2 Future Directions
Another promising area is multi-modal learning, which combines evidence from text with other modalities (e.g., cardiology images) to further enhance model performance in tasks such as the diagnosis of cardiac diseases. Building upon state-of-the-art multi-modal LLMs [222], there are emerging multi-modal methods that have shown promising results in information from ECG reports and images [203], visual QA from ECG image interpretations [223]. In future work, there is significant potential to combine further combine ECG or cardiac MRI images with narrative clinical text as more reliable information sources to improve model performance on disease diagnosis and risk prediction tasks.
@@
Previous open-source NLP tools like the GATE-based NLP system, I2E software, and cTAKES with UMLS CUIs are widely used in tasks like outpatient worsening heart failure identification [224] and information extraction from echocardiography reports [225]. These studies proved the usefulness and reliability of open-source NLP tools across health systems and use cases. Further development of LLM-based methods in the cardiology field can also significantly benefit from more powerful open-source LLM series for public use and high-quality open-source training data that corresponds to data privacy and related regulations. Overall, future research on open-source tools bears huge potential in democratizing AI usage in cardiology.

### v2

@@ Abstract
Cardiovascular diseases are becoming increasingly prevalent in modern society, with a profound impact on global health and well-being. These Cardiovascular disorders are complex and multifactorial, influenced by genetic predispositions, lifestyle choices, and diverse socioeconomic and clinical factors. Information about these interrelated factors is dispersed across multiple types of textual data, including patient narratives, medical records, and scientific literature. Natural language processing (NLP) has emerged as a powerful approach for analysing such unstructured data, enabling healthcare professionals and researchers to gain deeper insights that may transform the diagnosis, treatment, and prevention of cardiac disorders. This review provides a comprehensive overview of NLP research in cardiology from 2014 to 2025. We systematically searched six literature databases for studies describing NLP applications across a range of cardiovascular diseases. After a rigorous screening process, we identified 265 relevant articles. Each study was analysed across multiple dimensions, including NLP paradigms, cardiology-related tasks, disease types, and data sources. Our findings reveal substantial diversity within these dimensions, reflecting the breadth and evolution of NLP research in cardiology. A temporal analysis further highlights methodological trends, showing a progression from rule-based systems to large language models. Finally, we discuss key challenges and future directions, such as developing interpretable LLMs and integrating multimodal data. To the best of our knowledge, this review represents the most comprehensive synthesis of NLP research in cardiology to date.
@@ 1 Introduction
Cardiovascular diseases (CVDs) comprise a group of disorders that affect the heart and blood vessels, posing a major global health challenge. The most common clinical manifestations include coronary artery disease, heart failure, arrhythmias, valvular heart disease, and hypertension. These conditions can significantly diminish an individual’s quality of life and overall well-being, with their prevalence continuing to rise due to aging populations and lifestyle changes [1]. According to the World Health Organization, CVDs are the leading cause of death worldwide, accounting for an estimated 17.9 million deaths each year [2]. Consequently, early detection and effective management of CVDs are essential to improving patient outcomes and mitigating their burden on healthcare systems.
@@ 2.1 Search strategy
Table 1: The keywords for literature search.
@@ 3 Findings
  • The type of NLP methods employed;
  • The cardiology-related tasks/applications addressed by the research;
  • The type of cardiovascular diseases that forms the focus of the research;
  • The type of datasets to which the reported NLP method was applied.
@@ 3.1 Data Sources
Figure 3: The hierarchical dataset classification schema for classification of data sources.
@@ 3.1.1 Top-level Category
Figure 4: Pie chart of the distribution of data source categories in Cardiology NLP research.
@@
As shown in figure 4, a total of 9 different types of textual data have been considered in the reviewed articles. The vast majority of studies make use of Electronic Health Records (EHRs), which account for 82.4% of the data sources used in the studies included in the review. Published literature and medical databases constitute the second largest category at 4.7%, followed closely by synthetic or generated data at 4.3%. Clinical Trial Data represents 2.7% of the sources, while telehealth data accounts for 1.9%. Patient-reported outcomes and surveys make up 1.6%, and home healthcare data represents 1.2%. Administrative data (0.8%) and social media data (0.4%) represent smaller portions of the data sources used.
@@
Synthetic or Generated Data
@@
Clinical trial data represents highly structured and carefully collected information from controlled research studies investigating medical interventions, treatments, or diagnostic approaches [27, 25]. This data type is characterized by its rigorous collection protocols, standardized formats, and detailed documentation of patient outcomes and adverse events. Clinical trial data is considered the gold standard for evidence-based medicine, though its generalizability may be limited by strict inclusion/exclusion criteria. For example, Cunningham et al. [28] utilized the INVESTED trial data [29] to validate and fine-tune a Natural Language Processing (NLP) model for automatically identifying heart failure hospitalizations.
@@
Patient-reported Outcomes and Surveys
Home Healthcare Data
Home healthcare data encompasses clinical information collected during healthcare services provided in patients’ homes, including vital signs, symptoms, medication adherence, and functional assessments [35]. This data type offers unique insights into patients’ daily living conditions and health management behaviours outside traditional clinical settings. Home healthcare data is particularly valuable for understanding chronic disease management and patient self-care capabilities. For example, Chae [36] used 2.3 million home healthcare clinical notes to identify patients in home healthcare with heart failure symptoms and poor self-management.
Administrative data
Administrative data encompasses structured information generated during healthcare business operations, including insurance claims, billing records, and operational metrics, which is typically maintained in systems separate from but complementary to EHRs [37, 38]. While primarily collected for billing, reimbursement, and administrative purposes, source allocation, and system-level performance indicators across different healthcare settings, administrative data has become a valuable resource for healthcare research due to its comprehensiveness and longitudinal nature. In the reviewed papers, Sharma [39] utilised administrative health data to predict the risk of 30-day readmissions in patients with heart failure.
@@
Figure 5: Distribution of different subcategories of EHRs.
@@ 3.1.3 Clinical Documentation and Diagnostic Information
Figure 6: Distribution of sub-categories in clinical documentations and diagnostic information.
@@ 3.1.4 Imaging and Radiology Reports
Figure 7: Distribution of sub-categories in cardiac imaging and radiology reports.
@@
The co-occurrence of cardiac imaging and radiology reports reveals notable patterns in how these report types are used together in research. Echocardiogram reports frequently appear alongside chest X-ray and ECG reports, reflecting common research approaches that combine anatomical imaging with functional assessment. Cardiac MRI and CT reports often co-occur, suggesting their complementary roles as advanced imaging modalities. Coronary angiography reports also frequently appear with echocardiogram, ECG, and cardiac catheterization reports, highlighting the integrated, multimodal strategies commonly employed in cardiovascular research.
@@ 3.3 Target Cardiovascular Diseases
In Coronary Artery Disease research, NLP has been particularly valuable in analyzing cardiac catheterization reports [95, 96, 97] and identifying the risk of acute coronary events from clinical documentation [4, 98, 99]. Studies in this category have also focused on developing predictive models for disease onset and outcome assessment [100].
@@ 3.4.1 Rule-Based Methods
Rule-based methods use predefined patterns, dictionaries, and expert-crafted rules to extract information from text. These methods have proven particularly effective for extracting standardized measurements and well-defined clinical entities. For example, Wagholikar [112] extracted ejection fraction values from echocardiogram reports, which achieved high precision due to the consistent formatting of these measurements. Ambrosy [21] utilized rule-based approaches to leverage standardized clinical criteria for identifying worsening heart failure events. Olivella et al. [113] used rule-based NLP approaches to filter cases of heart failure patients. Mefford et al. [114] used rule-based NLP methods to examine variation in heart failure hospitalizations, considering multiple factors. There are also other works using rule-based statistics (e.g., Bag-of-Words, Skip-gram) as features for machine learning algorithms, successfully applied in scenarios such as random forest [115, 116] and support vector machine [117, 85, 118].
@@ 3.4.3 Deep Learning-based Methods
Since the release of ChatGPT in 2022, Large Language Models (LLMs) have marked a major paradigm shift in NLP. LLMs are huge Transformer-based models (usually billions of parameters) pre-trained on massive scale of web data (usually trillions of tokens) to obtain outstanding general problem-solving capabilities and strong in-context learning capabilities, which enable LLMs to adapt to new tasks without modifying the parameters. LLMs have become an emerging dominant paradigm in cardiology, significantly outperforming previous methods in risk prediction for cardiovascular diseases [181, 82, 150], heart failure [183], and automated echocardiography report analysis [182]. Many works also explored utilizing LLMs for assisting diagnosis of related diseases, including as knowledge sources [200], agent for symptom description and management [201], and interpreters of existing reports [202]. LLMs also show generalized ability to handle signals from multiple sources, including ECG signals [71, 203, 204], text-based cardiology disease reports [205, 206, 207], and educational materials [75, 87, 208, 209]. Due to their auto-regressive nature, LLMs are also widely used in text generation tasks, successfully applied in question answering for cardiology-related topics [24, 63, 184, 185], AI assistants for clinical advice on myocardial infarction [186, 187], atrial fibrillation [78, 188], cardiovascular rehabilitation [189], and cardiac electrophysiology procedures [210, 7], and writing summaries of cardiology reports [66], cardiovascular magnetic resonance reports [190, 211], and education materials [191], etc. LLMs are also explored as research and decision-support tools for domain experts [192, 212, 213]. In practical scenarios, the high-quality LLM-generated content can significantly lower the burden of domain experts and the cost of patients for professional advice. Another application focuses on extracting information from texts via prompt-based instructions. For example, Xu et al. [72] leveraged LLMs to extract structured knowledge from heart failure-related texts and built the first knowledge graph in this domain. Shang et al. [193] evaluated the factuality of extracted Cardiovascular-Kidney-Metabolic syndrome knowledge within LLMs to ensure their reliability. Chi et al. [194] utilized LLMs to extract echocardiogram entities in a flexible manner. Tully et al. [195] compared LLMs to rank the priorities for surgical patients needing preoperative cardiac evaluations. Vasisht Shankar et al. [214] converted text-based cardiovascular diagnostic reports into structured datasets with LLMs. Finally, due to LLMs’ high hallucination rate in knowledge-intensive scenarios, the retrieval-augmented generation (RAG) technique has been explored in cardiology, which aims to enhance LLM-generated content with retrieved knowledge or texts. For example, Adejumo et al. [81] incorporated structured information from inpatient heart failure notes to combine into the context of LLMs. Hayama et al. [196] comprehensively evaluated RAG-based LLMs and general LLMs in Cardiology specialist examinations in Japan, proving external knowledge useful for QA tasks. Lakhdhar et al. [197] used RAG in developing AI assistants for personalized cardiovascular risk assessment, and Parameswaran et al. [198] further proved its effectiveness by using RAG to enhance LLM performance in providing nutrition information for cardiovascular diseases.
@@ 4.1 Main Challenges
Data Privacy and Compliance with Regulations
@@ 4.2 Future Directions
Another promising direction is multi-modal learning, which combines evidence from text with other modalities to further enhance model performance in cardiology tasks. Building upon state-of-the-art multi-modal LLMs [223], there are emerging multi-modal methods that have shown promising results in integrating information from ECG reports and images [204], visual QA from ECG image interpretations [224]. In future work, there is significant potential to combine other cardiology images, such as cardiac MRI images, with narrative clinical text. Other modalities also bear potential as reliable information sources, including knowledge graphs [72, 193], other ECG signals [69], and heart rate variation signals [225], which can be leveraged to further improve model performance on disease diagnosis and risk prediction tasks.
@@
Previous open-source NLP tools like the GATE-based NLP system, I2E software, and cTAKES with UMLS CUIs are widely used in tasks like outpatient worsening heart failure identification [226] and information extraction from echocardiography reports [227]. These studies proved the usefulness and reliability of open-source NLP tools across health systems and use cases. Further development of LLM-based methods in the cardiology field can also significantly benefit from more powerful open-source LLM series for public use and high-quality open-source training data that corresponds to data privacy and related regulations. More open-source datasets for low-resource languages can also significantly benefit future research on NLP applications in multi-lingual scenarios. Overall, future research on open-source tools bears huge potential in democratizing AI usage in cardiology.
@@ 5 Conclusion
Acknowledgments
This work is supported by British Heart Foundation Manchester Centre of Research Excellence (RE/24/130017). This work is supported in part by The University of Manchester President’s Doctoral Scholar award.